OpenStudy (anonymous):

Chance can wash the car in 2 hours, while Sandy can wash the same car in 4 hours. How many hours would it take for them to wash the car together? four thirds three fourths 3 4

OpenStudy (anonymous):

Think you can help me @ganeshie8

OpenStudy (anonymous):

yes no maybe @ganeshie8

ganeshie8 (ganeshie8):

thinking lol.. this prob is spinning my head -.- gimme time to think a bit :)

ganeshie8 (ganeshie8):

Chance can wash the car in 2 hours that means he can complete 1/2 of work in 1 hour

ganeshie8 (ganeshie8):

Sandy can wash the same car in 4 hours. that means he can complete 1/4 of work in 1 hour

ganeshie8 (ganeshie8):

say, both together completes the work in \(x\) hours

ganeshie8 (ganeshie8):

then, \(\large \frac{1}{2} + \frac{1}{4} = \frac{1}{x}\)

ganeshie8 (ganeshie8):

solve \(x\)

ganeshie8 (ganeshie8):

see if u can solve \(x\) ...

OpenStudy (anonymous):

what i did was multiply 1*2/2*2 = 2/4 then I added that with 1/4 and I got 3/4 @ganeshie8

ganeshie8 (ganeshie8):

\(\large \frac{3}{4} = \frac{1}{x}\) \(\large x = \frac{4}{3} \)